OpenStudy (anonymous):

Find the value of n such that each expression is a perfect square trinomial. p^2+10p+n

myininaya (myininaya):

\[x^2+bx+n=x^2+bx+(\frac{b}{2})^2=(x+\frac{b}{2})^2 => n=(\frac{b}{2})^2\]